1. a small bathroom has two very hard walls on opposite sides of the room at a distance of 2.55m.  these walls reflect sound so that sound waves have nodes at the walls.  for which approximate frequency can a standing wave occur between the walls (resonance).


2. a container holding 1 liter is filled with air of room temperature (20 degrees Celsius) and normal atmospheric pressure.  it is compressed to half volume and cooled to -53 celsius, what is the pressure in the air.


4. an electron is accelerated from the rest of a static electric field by a potential difference of 0.51 MV.  what speed do the electrons get?


5. a point particle moves in space so that its position vector is given by: ((x (t), y (t), z (t)) = (f (t / x), g (2t / x), h (  3t / x)) where f, g and h are three periodic functions with period 1 and x is a constant, what is the average of a velocity vector from time t to t + x.


1
Expert's answer
2021-04-19T17:10:55-0400
  1. Given distance between the wall"(L)=\frac{n\lambda}{2}"

"\lambda = 2L =2\times 2.55 m= 5.10m"


2.


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"

Now, substituting the values,

"P_2=\frac{P_1V_1\times T_2}{T_1\times V_2}"


"=\frac{1\times 10^5\times 1\times 10^{-3}\times (273-53)\times 2}{(273+20)\times 10^{-3}}"

"=1.5\times 10^5 atm"


4.

Now, applying the conservation of energy,

"\frac{1}{2} mv^2=eV"

"\Rightarrow v=\sqrt{\frac{2eV}{m}}"

"\Rightarrow v = \sqrt{\frac{2\times 1.6\times 10^{-19}\times 0.51\times 10^{6}}{9.1\times 10^{-31}}} m/s"


"\Rightarrow v =\sqrt{0.18\times 10^{18}}m/s"


"\Rightarrow v = 0.424\times 10^{9}m/s"


5.


"x(t)=f(\frac{t}{x})\hat{i}"


"y(t)= g(\frac{2t}{x})\hat{j}"


"z(t)=h(\frac{3t}{x})\hat{k}"

"\overrightarrow{r}=x(t)+y(t)+z(t)"


"=f(\frac{t}{x})\hat{i}+g(\frac{2t}{x})\hat{j}+h(\frac{3t}{x})\hat{k}"


"v_{avg}=\frac{d\overrightarrow{r}}{dt}"


"=\frac{1}{x}f(\frac{t}{x})\hat{i}+\frac{2}{x}g(\frac{2t}{x})\hat{j}+\frac{3}{x}h(\frac{3t}{x})\hat{k}"